OpenStudy (anonymous):

Which of the following is true about the function below? 1/√x+4

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

its domains is \[(-\infty,0]\] and its range is \[(0,\infty)\] its domains is \[(-4,\infty)\] and its range is \[(0,\infty)\] its domains is \[(-4,\infty)\] and its range is \[(\infty,-\infty)\] its domains is \[[-4,\infty)\] and its range is \[(-\infty,0) \cup (0,\infty)\]

jhonyy9 (jhonyy9):

so like a first step hope you know when a fraction will be undefined ,yes ? when the denominator is zero so in this case when for what value of x will result the denominator equal zero ???